Q[x]/<x-5> isomorphic to Q as field.
1
Expert's answer
2020-12-17T08:20:22-0500

The elements of the field "\mathbb Q[x]/\langle x-5\rangle" have the form "[f(x)]=f(x)+(x-5)\mathbb Q[x]". According to the Polynomial Euclidean Algorithm for any polynomial "f(x)\in\mathbb Q[x]" there is a unique polinomial "r(x)\in \mathbb Q[x]" such that "f(x)=(x-5)q(x)+r(x),\ \deg(r(x))<\deg(x-5)=1," for some unidue polinomial "q(x)\in\mathbb Q[x]". It follows from "\deg(r(x))<1" that "r(x)=r\in\mathbb Q", and thus "[f(x)]=[r]." Therefore, "\mathbb Q[x]/\langle x-5\rangle=\{[r]\ :\ r\in\mathbb Q\}".

Let us define a map "f:\mathbb Q\to\mathbb Q[x]/\langle x-5\rangle, \ \ f(r)=[r]." Taking into account that "f(a+b)=[a+b]=[a]+[b]=f(a)+f(b)" and "f(a\cdot b)=[a\cdot b]=[a]\cdot [b]=f(a)\cdot f(b)", we conclude that "f" is a field homomorphism. Since the reminder is a unique, "a\ne b" implies "[a]\ne [b]", and therefore, "f" is injective. For each "[r]\in\mathbb Q[x]/\langle x-5\rangle" we have that "f(r)=[r]", and "f" is surjective. Consequently, "f" is a field isomorphism, and "\mathbb Q[x]/\langle x-5\rangle" isomorphic to "\mathbb Q" as field.